OpenStudy (anonymous):

A survey about favoritecolors reports that EXACTLY: 12% of those surveyed prefer red 14% of those surveyed prefer Orange 28% of those surveyed prefer purple 30% of those surveyed prefer Blue 16% of those surveyed prefer green What is the smallest number of people that could have taken the survey? Explain Updated 3 minutes ago

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

...

OpenStudy (anonymous):

so we want an integer , a positive integer solution

OpenStudy (anonymous):

I would start by expressing each as a fraction over 100 and reducing to get closer to finding the least common multiple of the denominators

OpenStudy (anonymous):

so total must be divisible by

OpenStudy (anonymous):

ENGLISH PEOPLE ENGLISH

OpenStudy (anonymous):

tons is right

OpenStudy (anonymous):

tonks*

OpenStudy (anonymous):

is anybody good at java programming

OpenStudy (anonymous):

In english what he said means?

OpenStudy (anonymous):

12/100= 3/25

OpenStudy (anonymous):

12 out of 100 is 6 out of 50 or 3 out of 25

OpenStudy (anonymous):

WHERE R U GETTING THESE NUMBERS

OpenStudy (anonymous):

while 14 out of 100 is 7 out of 50 and can't be reduced any further

OpenStudy (anonymous):

percent means out of 100

OpenStudy (anonymous):

...so 12% = 12 out of 100 or 12/100 which reduces like a fraction to 6/50 and 3/25

OpenStudy (anonymous):

28/100=7/25

OpenStudy (anonymous):

.30=3/10

OpenStudy (anonymous):

ya so the minimum ammount of people is 10?

OpenStudy (anonymous):

.16=4/25

OpenStudy (anonymous):

ok so we want a number that is divisible by all those denominators

OpenStudy (anonymous):

Im not good with m,ath but i can rock your orld with any other subject >.<

OpenStudy (anonymous):

So, if you look at all of them, they all can be out of 100, obviously, but there is a smaller number that they all can be out of

OpenStudy (anonymous):

can they all be expressed as a whole number out of 25 by reducing like in the process above?

OpenStudy (anonymous):

Uhhh

OpenStudy (anonymous):

or rather

OpenStudy (anonymous):

idk...?

OpenStudy (anonymous):

like 12% = 12/100 = 3/25, can they all go to out of 25?

OpenStudy (anonymous):

multiply all the denominators

OpenStudy (anonymous):

SO MANY WORDS... its like a mathbook

OpenStudy (anonymous):

im not sure if they can

OpenStudy (anonymous):

I am using "like" and "out of", do you want help or the answer

OpenStudy (anonymous):

Help What do you mean by "Like"

OpenStudy (anonymous):

for instance

OpenStudy (anonymous):

so the smallest number sampled possibly is 25*50*25*10*25

OpenStudy (anonymous):

We're not on facebook.... lol

OpenStudy (anonymous):

tonks, do you concur

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

no, I think there is a faster way, by noticing what they can reduce to

OpenStudy (anonymous):

I DONT SPEAK MATH.... MUST.... FIND.... TANSLATOR.....

OpenStudy (anonymous):

i think this is the only way

OpenStudy (anonymous):

reduce the fractions to lowest terms, and then find the product of the denominators?

OpenStudy (anonymous):

oh theres redundancy

OpenStudy (anonymous):

WHAT IS A LOWEST TERM?

OpenStudy (anonymous):

like 3/6 = 1/2 , lowest terms

OpenStudy (anonymous):

and what's redundancy... IM IN THE 7TH GRADE

OpenStudy (anonymous):

You mean shrink them....

OpenStudy (anonymous):

so the answer is 50

OpenStudy (anonymous):

12% of those surveyed prefer red = 12/100 = 6/50 = 3/25 14% of those surveyed prefer Orange = 7/50 28% of those surveyed prefer purple = 28/100 = 14/50 30% of those surveyed prefer Blue = 30/100 = 15/50 16% of those surveyed prefer green = 16/100 = 8/50 Notice that MOST of them can be expressed as a whole number out of 25, but not ALL of them...

OpenStudy (anonymous):

all the denominators go into that, thats the lowest common multipled of 25,50,25,10,25

OpenStudy (anonymous):

so then its 50

OpenStudy (anonymous):

While ALL of them can be rewritten as a whole number out of 50, so the answer is....

OpenStudy (anonymous):

so how do i figure it out? xD

OpenStudy (anonymous):

you rub a magic lamp, and say , genie genie please give me answer

OpenStudy (anonymous):

^^^^^

OpenStudy (anonymous):

if that doesnt work, you can always pay me

OpenStudy (anonymous):

Cantor ill give you a medal

OpenStudy (anonymous):

MY BRAIN HURTS D:

OpenStudy (anonymous):

12% of those surveyed prefer red = 12/100 = 6/50 = 3/25 14% of those surveyed prefer Orange = 7/50 28% of those surveyed prefer purple = 28/100 = 14/50 30% of those surveyed prefer Blue = 30/100 = 15/50 16% of those surveyed prefer green = 16/100 = 8/50 Notice that MOST of them can be expressed as a whole number out of 25, but not ALL of them. While all of them can be rewritten as a fraction with a whole number on top and 50 in the bottom. The top of the fraction is the part and the bottom is the amount that it was "out of".

OpenStudy (anonymous):

Good luck, Zelando!

OpenStudy (anonymous):

so the smallest number of people is 50?

OpenStudy (anonymous):

ok zelando, say you have 12/100 , right ?

OpenStudy (anonymous):

YES!

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

12/100 * total must equal to an integer

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

I have smarticles!

OpenStudy (anonymous):

12/100 = 3/25, so we need a number that cancel with the denominator

OpenStudy (anonymous):

at the very least

OpenStudy (anonymous):

TOnks - you earned yourself a medal and you too can :)

OpenStudy (anonymous):

so like 12/100 * 25 works, since you get 12/100 = 3/25 and 3/25 * 25 = 3

OpenStudy (anonymous):

Wow, thanks

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

ok so far?

OpenStudy (anonymous):

Can - its 50 lmao

OpenStudy (anonymous):

i know, im building up to that

OpenStudy (anonymous):

lol all drematic

OpenStudy (anonymous):

zelando, are you good at analytic geometry , i got a tough problem

OpenStudy (anonymous):

find an equation of two tangents through a point outside a circle to a given circle

OpenStudy (anonymous):

so youre given P(a,b), do you know calculus ?

OpenStudy (anonymous):

tonka, are we good?

OpenStudy (anonymous):

What's Analytic geometry?

OpenStudy (anonymous):

<---- 7th grader

OpenStudy (anonymous):

geometry plus xy plane

OpenStudy (anonymous):

^-----

OpenStudy (anonymous):

I know how to find dimater of a circle...

OpenStudy (anonymous):

oh nevermind

OpenStudy (anonymous):

lol